OpenStudy (anonymous):

Tiya earns $7 an hour mowing her neighbor's lawn. Part A: Create two variables and determine which is dependent and which is independent for this situation.

OpenStudy (anonymous):

Part A: n = number of hours spent mowing the lawn. 7 is the monehy per hour. i just need to know which is dependable and which is independent variable.

OpenStudy (anonymous):

@vera_ewing

OpenStudy (anonymous):

nvm.

OpenStudy (anonymous):

question closed i found it out myself xD

OpenStudy (mathstudent55):

The amount she earns is $7 for each hour she works. If she works 1 hour, she earns $7 * 1 = $7 If she works 2 hours, she earns $7 * 2 = $14 If she works 15 hours, she earns $7 * 15 = $105 Any number of hours you pick, you can find the amount of money she earns.

OpenStudy (mathstudent55):

n can be any number you want, so n is "independent." Let's say you name s as the salary. Then the equation to find her salary is s = 7n The salary is 7 times the number of hours. s always depends on n, so s is "dependent."